# A Tutorial on Beyond-Diagonal Reconfigurable Intelligent Surfaces: Modeling, Architectures, System Design and Optimization, and Applications

### v1

@@ V BD-RIS with Hardware Impairments
The discussions on BD-RIS so far have mainly focused on idealized and simplified hardware, such as lossless reconfigurable impedance networks with continuously tunable admittance components, with no frequency dependence, and with no mutual coupling between elements. These idealized assumptions help to understand the fundamentals and study the performance limit of BD-RIS. However, they cannot be perfectly achieved in real-world implementations. In this section, we thus briefly revisit some important hardware impairments of BD-RIS and their impacts on system performance.
V-A Discrete-Value Impedance and Admittance
@@
where $B$ denotes the number of resolution bits. In D-RIS literature, typical beamforming and channel estimation solutions are to quantize continuous-value results to (90), or to directly design each scattering coefficient by selecting from (90) [9]. This, however, does not work for BD-RIS with scattering matrices whose entries are dependent on each other. To model the discrete-value BD-RIS, [89] has focused on the lossless fully/group-connected architectures, and, for the first time, proposed to model the discrete-value reactance matrix $\mathbf{X}_{I}=\Im\{\mathbf{Z}_{I}\}=\mathsf{blkdiag}(\mathbf{X}_{I,1},\ldots,\mathbf{X}_{I,g})$, where $\mathbf{X}_{I,g}\in\mathbb{R}^{\bar{M}\times\bar{M}}$, $\mathbf{X}_{I,g}=\mathbf{X}_{I,g}^{\mathsf{T}}$ denotes the reactance matrix for group $g$ of the group-connected architecture. Then, each entry of $\mathbf{X}_{I,g}$ is constrained by
$[\mathbf{X}_{I,g}]_{m,m^{\prime}}\in\{\pm X_{I,b}~{}|~{}b\in\mathcal{B}\},\forall g\in\mathcal{G},\forall m,m^{\prime}\in\bar{\mathcal{M}},$ (91)
@@
where $X_{I,b}>0$, $\forall b\in\mathcal{B}$. Comparing (90) and (91), we observe that the codebook for D-RIS is known based on a finite interval, while that for BD-RIS contains unknown candidates ±$\pm X_{I,b}$. Therefore, the design of discrete-value BD-RIS includes an additional step, that is to determine the codebook first. To effectively solve this problem, [89] has proposed an offline learning method to obtain the codebook, which relies on a training set including sufficient channel realizations. With the determined codebook, the online deployment based on instantaneous CSI is conducted by iteratively designing individual reactance $[\mathbf{X}_{I,g}]_{m,m^{\prime}}$ through a one-dimensional exhaustive search. Based on the above design, results in Fig. 14 show that one resolution bit is sufficient in fully-connected BD-RIS to achieve satisfactory performance close to the continuous-value case. Such a mild requirement of the resolution bits is beneficial for the practical implementation of BD-RIS.
@@
Figure 14: Average received signal power versus the group size. The transmitter, RIS, and receiver are located at a 3D coordinate system with respectively (5,-250,25), (0,0,5), and (5,5,1.5) in meters (m). The channels through RIS have i.i.d. Rayleigh fading. ($N=4$, $N_{\mathsf{r}}=2$, $M=64$) [89].
@@
The design in [89] is also applicable to other reciprocal architectures. To have a unified framework of the discrete-value BD-RIS design, it is suggested to discretize the nonzero entries of admittance matrices as explained in Remarks 4 and 5. That is, given a lossless and reciprocal reconfigurable impedance network with admittance matrix $\mathbf{Y}_{I}$, each nonzero entry of the corresponding susceptance matrix $\mathbf{B}_{I}=\Im\{\mathbf{Y}_{I}\}\in\mathbb{R}^{M\times M}$ is constrained by
@@
$\forall~{}[\mathbf{B}_{I}]_{m,m^{\prime}}\neq 0,m,m^{\prime}\in\mathcal{M},$
where $B_{I,b}>0$, $\forall b\in\mathcal{B}$. The offline codebook design problem aiming at maximizing the average strength of a MIMO channel can be formulated as
@@
where $\mathbf{H}(\mathbf{\Theta})=\mathbf{H}_{RT}+\mathbf{H}_{RI}\mathbf{\Theta}\mathbf{H}_{IT}$ according to Section IV-B. The proposed offline learning method in [89] is readily adapted to BD-RIS architectures with different constraints $\mathcal{Y}$ as illustrated in Section III-C. Once the codebook design is obtained, the discrete-value BD-RIS can be designed based on iteratively exhaustive search as in [89], or by directly quantizing the continuous-value solutions from Section IV-B.
@@
There has not been existing literature solving exactly the problem (95), while it can be viewed as a vector quantization problem and potentially solved by the Lloyd algorithm [90]. The detailed design will be left as an interesting future work.
V-B Lossy Interconnections and Admittance Components
Power loss is a very important factor in hardware devices. Modeling the hardware without capturing accurately the loss could cause non-negligible performance degradation. In D-RIS, the power loss comes from the circuit of each reconfigurable component [17]. However, in BD-RIS architectures, the power loss comes from two aspects: 1) the interconnection loss between ports of the reconfigurable impedance network and 2) the loss from the circuit of reconfigurable admittance components. These two kinds of losses have been recently modeled and analyzed, respectively in [91] and [92], both focusing on reciprocal BD-RIS architectures. Below we will briefly revisit how to capture these two kinds of losses in BD-RIS architectures.
V-B.1 Lossy Interconnections
@@
The reciprocal BD-RIS with lossy interconnections is modeled as an $M$-port reconfigurable impedance network, characterized by its admittance matrix $\mathbf{Y}_{I}\in\mathbb{C}^{M\times M}$. Similar to the model in Section III-A, each port $m$ is connected to a reconfigurable admittance component $Y_{m}$ to ground, as illustrated in Fig. 15(a). The main difference is port $m$ and port $m^{\prime}$, $m^{\prime}\neq m$, if interconnected, are interconnected via a reconfigurable admittance component $Y_{m,m^{\prime}}$ in series with a transmission line with length $\ell_{m,m^{\prime}}$, $\forall m,m^{\prime}\in\mathcal{M}$, as illustrated in Fig. 15(b). According to the derivation in [91], the entries in $\mathbf{Y}_{I}$ have expressions
$[\mathbf{Y}_{I}]_{m,m^{\prime}}=\begin{cases}\frac{-2}{Y_{m,m^{\prime}}^{-1}\zeta_{m,m^{\prime}}^{+}+Z_{0}\zeta_{m,m^{\prime}}^{-}},&m\neq m^{\prime},\\ Y_{m}-\sum_{n\neq m}\frac{\zeta_{m,m^{\prime}}^{+}}{2}[\mathbf{Y}_{I}]_{m,n},&m=m^{\prime},\end{cases}$ (96)
@@
where $\zeta_{m,m^{\prime}}^{+}=e^{\zeta\ell_{m,m^{\prime}}}+e^{-\zeta\ell_{m,m^{\prime}}}$ and $\beta_{m,m^{\prime}}^{-}=e^{\zeta\ell_{m,m^{\prime}}}-e^{-\zeta\ell_{m,m^{\prime}}}$ with $\zeta\in\mathbb{C}$ being the propagation constant of the transmission line. Specifically, in $\zeta=\alpha+\jmath\beta$, the real part $\alpha$ denotes the attenuation constant and the imaginary part $\beta$ denotes the phase constant. It is difficult to gain insights from (96) given its complex expressions. To understand how the lossy interconnections in BD-RIS impact the feasible range of entries in $\mathbf{Y}_{I}$, [91] has simplified (96) by assuming a special case where $\ell_{m,m^{\prime}}$ is a multiple of half of wavelength $\lambda$, $\lambda=\frac{2\pi}{\beta}$. That is, $\ell_{m,m^{\prime}}=\frac{\pi}{\beta}A_{m,m^{\prime}}$, where $A_{m,m^{\prime}}\in\mathbb{Z}$. In addition, to focus purely on the impact of lossy interconnections, each reconfigurable admittance component is assumed to be lossless, i.e., $\Re\{Y_{m}\}=0$, $\Re\{Y_{m,m^{\prime}}\}=0$, $\forall m,m^{\prime}\in\mathcal{M}$. Then, the constraint of $[\mathbf{Y}_{I}]_{m,m^{\prime}}$ with $m\neq m^{\prime}$ can be simplified to
$[\mathbf{Y}_{I}]_{m,m^{\prime}}=\frac{-(-1)^{A_{m,m^{\prime}}}}{\underbrace{Y_{m,m^{\prime}}^{-1}\cosh(\alpha\ell_{m,m^{\prime}})}_{\text{purely imaginary}}+\underbrace{Z_{0}\sinh(\alpha\ell_{m,m^{\prime}})}_{\text{purely real}}},$ (97)
@@
indicating that $[\mathbf{Y}_{I}]_{m,m^{\prime}}$, if is not zero, must have a non-zero real part resulting in power loss. More interestingly, the existence of lossy interconnections makes the value of $[\mathbf{Y}_{I}]_{m,m^{\prime}}$ being constrained on a circle in the complex plane. An extreme case ($\alpha=0$) that the real parts of $[\mathbf{Y}_{I}]_{m,m^{\prime}}$ are zeros corresponds to BD-RIS with lossless interconnections. The constraint (97) referring to BD-RIS with lossy interconnections and lossless BD-RIS can be visualized in Fig. 16. From Fig. 16 we observe that the real and imaginary parts of $[\mathbf{Y}_{I}]_{m,m^{\prime}}$ are dependent on each other in lossy cases and the possible value of $[\mathbf{Y}_{I}]_{m,m^{\prime}}$ depends on the value of $\alpha$, while the imaginary part of $[\mathbf{Y}_{I}]_{m,m^{\prime}}$ for lossless cases can take arbitrary values, consistent with existing BD-RIS literature [26, 40, 57]. The impact of lossy interconnections in BD-RIS to system performance has also been shown in Fig. 17, from which we observe that the impact of lossy interconnections in BD-RIS is negligible due to the small inter-element spacing.
@@
Figure 16: Values of $[\mathbf{Y}_{I}]_{m,m^{\prime}}$ modeled as (97) with $Z_{0}=50\Omega$ and $\ell_{m,m^{\prime}}=0.1$ m [91].
@@
Figure 17: Average received signal power versus the number of RIS elements with different values of $\alpha$. The transmitter, RIS, and, receiver are located at a 3D coordinate system with respectively (0,0,0), (20,0,2), and (20,0,0) in meters (m). The inter-element spacing in BD-RIS is set as 0.05 m. Channels through RIS follow i.i.d. Rayleigh fading ($N=N_{\mathsf{r}}=1$) [91].
@@
V-B.2 Lossy Admittance Components
In D-RIS, the power loss of each reconfigurable component can be individually characterized in the amplitude of its reflection coefficients [17]. That is, for a lossy D-RIS, each entry $\varTheta_{m}$ in the scattering matrix has $|\varTheta_{m}|<1$. The smaller the value of $|\varTheta_{m}|$, the larger the loss in D-RIS. However, it is not possible to model individually the entries in the scattering matrix $\mathbf{\Theta}$ of BD-RIS since they are coupled with each other due to interconnections. To tackle this difficulty, [92] has recently proposed to directly analyze the loss of each reconfigurable admittance component, and the scattering matrix for lossy BD-RIS can be naturally obtained by (17) and (23). This is done by modeling each reconfigurable admittance component as a lumped circuit consisting of inductors $L_{1}$, $L_{2}$, a tunable capacitance $C$, and an equivalent resistor $R$ that accounts for the parasitic resistance of the varactor, as illustrated in Fig. 18. The admittance (if is not zero), with a given frequency of signals $\omega_{\mathsf{c}}=2\pi f_{\mathsf{c}}$, is a function of the tunable capacitance $C$ given by
@@ V-B.2 Lossy Admittance Components
($\Big{(}\Re\{Y_{m,m^{\prime}}\}-\frac{1}{2R}\Big{)}^{2}+\Big{(}\Im\{Y_{m,m^{\prime}}\}-\Big{(}-\frac{1}{\omega_{\mathsf{c}}L_{1}}\Big{)}\Big{)}^{2}=\Big{(}\frac{1}{2R}\Big{)}^{2}.$ (99)
@@
Figure 19: Values of $Y_{m,m^{\prime}}$ modeled as (98). The frequency is set as $f_{\mathsf{c}}=2.4$ GHz. $L_{1}=6$ nH, $L_{2}=0.7$ nH. Circles represent all possible values of $Y_{m,m^{\prime}}$ for varing $R$, while the practical range is constrained by tha value of capacitance $C\in[0.35,3.2]$ pF [92].
@@
Figure 20: Average rate versus $R$ for a lossy BD-RIS-aided SISO system. The channels through BD-RIS follow Rician fading with a Rician factor 2 dB, and the direct channel from the transmitter to the receiver follows Rayleigh fading. The transmitter-receiver, transmitter-RIS, and RIS-receiver distances are respectively set as 52 m, 50 m, and 2.5 m ($M=30$, $\bar{M}=6$ for the group/forest-connected architecture, $P=20$ dBm) [92].
@@
The above modeling of lossy interconnections and admittance components is applicable to all reciprocal BD-RIS architectures as illustrated in Section III-C. Given a specific circuit topology, one can simply set $Y_{m,m^{\prime}}=0$ to indicate that there are no direct interconnections between port $m$ and $m^{\prime}$, and adopt (96) or (97) to construct the admittance matrix with lossy interconnections, and to adopt (98) and (23) to construct the admittance matrix with lossy reconfigurable admittance components. However, the joint consideration of the impact of two kinds of losses to system performance remains unexplored and is a meaningful future research direction.
V-C Wideband Effect

### v2

@@ VI BD-RIS with Hardware Impairments
The discussions on BD-RIS so far have mainly focused on idealized and simplified hardware, such as lossless reconfigurable impedance networks with continuously tunable admittance components, with no frequency dependence, and with no mutual coupling between elements. These idealized assumptions help to understand the fundamentals and study the performance limit of BD-RIS. However, they cannot be perfectly achieved in real-world implementations. In this section, we thus briefly revisit some important hardware impairments of BD-RIS and their impacts on system performance. Similar to Section IV-C where we discussed channel estimation errors alone assuming perfect hardware at BD-RIS, below we will also show the impact of each hardware impairment assuming perfect other factors to guide readers from scratch. Analyzing real-world BD-RIS-aided scenarios by jointly considering multiple practical factors can be a meaningful future direction.
VI-A Discrete-Value Impedance and Admittance
@@
where $B$ denotes the number of resolution bits. In D-RIS literature, typical beamforming and channel estimation solutions are to quantize continuous-value results to (88), or to directly design each scattering coefficient by selecting from (88). This, however, does not work for BD-RIS with scattering matrices whose entries are dependent on each other. To model the discrete-value BD-RIS, [nerini2023discrete] has focused on the lossless fully/group-connected architectures, and, for the first time, proposed to model the discrete-value reactance matrix $\mathbf{X}_{I}=\Im\{\mathbf{Z}_{I}\}=\mathsf{blkdiag}(\mathbf{X}_{I,1},\ldots,\mathbf{X}_{I,g})$, where $\mathbf{X}_{I,g}\in\mathbb{R}^{\bar{M}\times\bar{M}}$, $\mathbf{X}_{I,g}=\mathbf{X}_{I,g}^{\mathsf{T}}$ denotes the reactance matrix for group $g$ of the group-connected architecture. Then, each entry of $\mathbf{X}_{I,g}$ is constrained by [^6] In D-RIS, the reactance matrix $\mathbf{X}_{I}$ boils down to a diagonal matrix and each diagonal entry $\mathbf{X}_{I}]_{m,m}$ is linked to $\varTheta_{m}$ in $\mathbf{\Theta}$ as $\varTheta_{m}=\frac{\jmath[\mathbf{X}_{I}]_{m,m}+Z_{0}}{\jmath[\mathbf{X}_{I}]_{m,m}-Z_{0}}$, such that one can directly obtain the discrete-value phase shift $\angle\varTheta_{m}$ with a discrete-value reactance $\mathbf{X}_{I}]_{m,m}$.
$\mathbf{X}_{I,g}]_{m,m^{\prime}}\in\{\pm X_{I,b}~|~b\in\mathcal{B}\},\forall g\in\mathcal{G},\forall m,m^{\prime}\in\bar{\mathcal{M}},$ (89)
@@
where $X_{I,b}>0$, $\forall b\in\mathcal{B}$. Comparing (88) and (89), we observe that the codebook for D-RIS is known based on a finite interval, while that for BD-RIS contains unknown candidates $\pm X_{I,b}$. Therefore, the design of discrete-value BD-RIS includes an additional step, that is to determine the codebook first. To effectively solve this problem, [nerini2023discrete] has proposed an offline learning method to obtain the codebook, which relies on a training set including sufficient channel realizations. With the determined codebook, the online deployment based on instantaneous CSI is conducted by iteratively designing individual reactance $\mathbf{X}_{I,g}]_{m,m^{\prime}}$ through a one-dimensional search. Based on the above design, results in Fig. 25 show that one resolution bit is sufficient in fully-connected BD-RIS to achieve satisfactory performance close to the continuous-value case. Such a mild requirement of the resolution bits is beneficial for the practical implementation of BD-RIS.
@@
Figure 25: Average received signal power versus the group size. The transmitter, RIS, and receiver are located at a 3D coordinate system with respectively (5,-250,25), (0,0,5), and (5,5,1.5) in meters (m). The channels from the transmitter to receiver and through RIS have i.i.d. Rayleigh fading. ($N=4$, $N_{\mathsf{r}}=2$, $M=64$) [nerini2023discrete].
@@
The design in [nerini2023discrete] is also applicable to other reciprocal architectures. To have a unified framework of the discrete-value BD-RIS design, it is suggested to discretize the nonzero entries of admittance matrices as explained in Remarks 4 and 5. That is, given a lossless and reciprocal reconfigurable impedance network with admittance matrix $\mathbf{Y}_{I}$, each nonzero entry of the corresponding susceptance matrix $\mathbf{B}_{I}=\Im\{\mathbf{Y}_{I}\}\in\mathbb{R}^{M\times M}$ is constrained by
@@
for $\mathbf{B}_{I}]_{m,m^{\prime}}\neq 0,m,m^{\prime}\in\mathcal{M}$, where $B_{I,b}>0$, $\forall b\in\mathcal{B}$. The offline codebook design problem aiming at maximizing the average strength of a MIMO channel is formulated as
@@
where $\mathbf{H}(\mathbf{\Theta})=\mathbf{H}_{RT}+\mathbf{H}_{RI}\mathbf{\Theta}\mathbf{H}_{IT}$ according to Section IV-B. The proposed offline learning method in [nerini2023discrete] is readily adapted to BD-RIS architectures with different constraints $\mathcal{Y}$ as illustrated in Section III-D. Once the codebook design is obtained, the discrete-value BD-RIS can be designed based on iteratively exhaustive search as in [nerini2023discrete], or by directly quantizing the continuous-value solutions from Section IV-B.
@@
There has not been existing literature solving exactly the problem (93), while it can be viewed as a vector quantization problem and potentially solved by the Lloyd algorithm [sabin1986global]. The detailed design will be left as an interesting future work.
VI-B Lossy Interconnections and Admittance Components
Power loss is a very important factor in hardware devices. Modeling the hardware without capturing accurately the loss could cause non-negligible performance degradation. In D-RIS, the power loss comes from the circuit of each reconfigurable component [abeywickrama2020intelligent]. However, in BD-RIS architectures, the power loss comes from two aspects: 1) the interconnection loss between ports of the reconfigurable impedance network and 2) the loss from the circuit of reconfigurable admittance components. These two kinds of losses have been recently modeled and analyzed, respectively in [nerini2025localized] and [peng2025lossy], both focusing on reciprocal BD-RIS architectures. Below we will briefly revisit how to capture these two kinds of losses in BD-RIS architectures.
VI-B.1 Lossy Interconnections
@@
The reciprocal BD-RIS with lossy interconnections is modeled as an $M$-port reconfigurable impedance network, characterized by its admittance matrix $\mathbf{Y}_{I}\in\mathbb{C}^{M\times M}$. Similar to the model in Section III-B, each port $m$ is connected to a reconfigurable admittance component $Y_{m}$ to ground, as illustrated in Fig. 26(a). The main difference is port $m$ and port $^{\prime}, mm$$^{\prime}\neq m$, if interconnected, are interconnected via a reconfigurable admittance component $Y_{m,m^{\prime}}$ in series with a transmission line with length $\ell_{m,m^{\prime}}$, $\forall m,m^{\prime}\in\mathcal{M}$, as illustrated in Fig. 26(b). According to the derivation in [nerini2025localized], the entries in $\mathbf{Y}_{I}$ have expressions
$\mathbf{Y}_{I}]_{m,m^{\prime}}=\begin{cases}\frac{-2}{Y_{m,m^{\prime}}^{-1}\zeta_{m,m^{\prime}}^{+}+Z_{0}\zeta_{m,m^{\prime}}^{-}},&m\neq m^{\prime},\\ Y_{m}-\sum_{n\neq m}\frac{\zeta_{m,m^{\prime}}^{+}}{2}[\mathbf{Y}_{I}]_{m,n},&m=m^{\prime},\end{cases}$ (94)
@@
where $\zeta_{m,m^{\prime}}^{+}=e^{\zeta\ell_{m,m^{\prime}}}+e^{-\zeta\ell_{m,m^{\prime}}}$ and $\beta_{m,m^{\prime}}^{-}=e^{\zeta\ell_{m,m^{\prime}}}-e^{-\zeta\ell_{m,m^{\prime}}}$ with $\zeta\in\mathbb{C}$ being the propagation constant of the transmission line. Specifically, in $\zeta=\alpha+\jmath\beta$, the real part $\alpha$ denotes the attenuation constant [^7] The value of $\alpha$ depends on the transmission line parameters, including series resistance and inductance, and shunt conductance and capacitance per unit length of the lumped-element equivalent circuit [pozar2021microwave]. and the imaginary part $\beta$ denotes the phase constant. It is difficult to gain insights from (94) given its complex expressions. To understand how the lossy interconnections in BD-RIS impact the feasible range of entries in $\mathbf{Y}_{I}$, [nerini2025localized] has simplified (94) by assuming a special case where $\ell_{m,m^{\prime}}$ is a multiple of half of wavelength $\lambda$, $\lambda=\frac{2\pi}{\beta}$. That is, $\ell_{m,m^{\prime}}=\frac{\pi}{\beta}A_{m,m^{\prime}}$, where $A_{m,m^{\prime}}\in\mathbb{Z}$. In addition, to focus purely on the impact of lossy interconnections, each reconfigurable admittance component is assumed to be lossless, i.e., $\Re\{Y_{m}\}=0$, $\Re\{Y_{m,m^{\prime}}\}=0$, $\forall m,m^{\prime}\in\mathcal{M}$. Then, the constraint of $\mathbf{Y}_{I}]_{m,m^{\prime}}$ with $m\neq m^{\prime}$ can be simplified to
[…]
@@
indicating that $\mathbf{Y}_{I}]_{m,m^{\prime}}$, if is not zero, must have a non-zero real part resulting in power loss. More interestingly, the existence of lossy interconnections makes the value of $\mathbf{Y}_{I}]_{m,m^{\prime}}$ being constrained on a circle in the complex plane. An extreme case ($\alpha=0$) that the real parts of $\mathbf{Y}_{I}]_{m,m^{\prime}}$ are zeros corresponds to BD-RIS with lossless interconnections. The constraint (95) referring to BD-RIS with lossy interconnections and lossless BD-RIS can be visualized in Fig. 27. From Fig. 27 we observe that the real and imaginary parts of $\mathbf{Y}_{I}]_{m,m^{\prime}}$ are dependent on each other in lossy cases and the possible value of $\mathbf{Y}_{I}]_{m,m^{\prime}}$ depends on the value of $\alpha$, while the imaginary part of $\mathbf{Y}_{I}]_{m,m^{\prime}}$ for lossless cases can take arbitrary values, consistent with existing BD-RIS literature [shen2021, nerini2024beyond, wu2024optimization]. The impact of lossy interconnections in BD-RIS to system performance has also been shown in Fig. 28, from which we observe that the impact of lossy interconnections in BD-RIS is negligible due to the small inter-element spacing.
@@
Figure 27: Values of $\mathbf{Y}_{I}]_{m,m^{\prime}}$ modeled as (95) with $Z_{0}=50\Omega$ and $\ell_{m,m^{\prime}}=0.1$ m [nerini2025localized].
@@
Figure 28: Average received signal power versus the number of RIS elements with different values of $\alpha$. The transmitter, RIS, and, receiver are located at a 3D coordinate system with respectively (0,0,0), (20,0,2), and (20,0,0) in meters (m). The inter-element spacing in BD-RIS is set as 0.05 m. The direct transmitter-receiver channel is assumed to be fully obstructed. Channels through RIS follow i.i.d. Rayleigh fading ($N=N_{\mathsf{r}}=1$) [nerini2025localized].
@@
VI-B.2 Lossy Admittance Components
In D-RIS, the power loss of each reconfigurable component can be individually characterized in the amplitude of its reflection coefficients [abeywickrama2020intelligent]. That is, for a lossy D-RIS, each entry $\varTheta_{m}$ in the scattering matrix has $|\varTheta_{m}|<1$. The smaller the value of $|\varTheta_{m}|$, the larger the loss in D-RIS. However, it is not possible to model individually the entries in the scattering matrix $\mathbf{\Theta}$ of BD-RIS since they are coupled with each other due to interconnections. To tackle this difficulty, [peng2025lossy] has recently proposed to directly analyze the loss of each reconfigurable admittance component, and the scattering matrix for lossy BD-RIS can be naturally obtained by (16) and (21). This is done by modeling each reconfigurable admittance component as a lumped circuit consisting of inductors $L_{1}$, $L_{2}$, a tunable capacitance $C$, and an equivalent resistor $R$ that accounts for the parasitic resistance of the varactor, as illustrated in Fig. 29. The admittance (if is not zero), with a given frequency of signals $\omega_{\mathsf{c}}=2\pi f_{\mathsf{c}}$, is a function of the tunable capacitance $C$ given by
@@ VI-B.2 Lossy Admittance Components
$\Big(\Re\{Y_{m,m^{\prime}}\}-\frac{1}{2R}\Big)^{2}+\Big(\Im\{Y_{m,m^{\prime}}\}-\Big(-\frac{1}{\omega_{\mathsf{c}}L_{1}}\Big)\Big)^{2}=\Big(\frac{1}{2R}\Big)^{2}.$ (97)
@@
Figure 30: Values of $Y_{m,m^{\prime}}$ modeled as (96). The frequency is set as $f_{\mathsf{c}}=2.4$ GHz. $L_{1}=6$ nH, $L_{2}=0.7$ nH. Circles represent all possible values of $Y_{m,m^{\prime}}$ for varing $R$, while the practical range is constrained by tha value of capacitance $C\in[0.35,3.2]$ pF [peng2025lossy].
@@
Figure 31: Average rate versus $R$ for a lossy BD-RIS-aided SISO system. The channels through BD-RIS follow Rician fading with a Rician factor 2 dB, and the direct channel from the transmitter to the receiver follows Rayleigh fading. The transmitter-receiver, transmitter-RIS, and RIS-receiver distances are respectively set as 52 m, 50 m, and 2.5 m ($M=30$, $\bar{M}=6$ for the group/forest-connected architecture, $P=20$ dBm) [peng2025lossy].
@@
The above modeling of lossy interconnections and admittance components is applicable to all reciprocal BD-RIS architectures as illustrated in Section III-D. Given a specific circuit topology, one can simply set $Y_{m,m^{\prime}}=0$ to indicate that there are no direct interconnections between port $m$ and $m^{\prime}$, and adopt (94) or (95) to construct the admittance matrix with lossy interconnections, and to adopt (96) and (21) to construct the admittance matrix with lossy reconfigurable admittance components. However, the joint consideration of the impact of two kinds of losses to system performance remains unexplored and is a meaningful future research direction.
VI-C Wideband Effect